# One Prompt Fits All: Universal Graph Adaptation for Pretrained Models

### v1

@@ Abstract
Graph Prompt Learning (GPL) has emerged as a promising paradigm that bridges graph pretraining models and downstream scenarios, mitigating label dependency and the misalignment between upstream pretraining and downstream tasks. Although existing GPL studies explore various prompt strategies, their effectiveness and underlying principles remain unclear. We identify two critical limitations: (1) Lack of consensus on underlying mechanisms: Despite current GPLs have advanced the field, there is no consensus on how prompts interact with pretrained models, as different strategies intervene at varying spaces within the model, i.e., input-level, layer-wise, and representation-level prompts. (2) Limited scenario adaptability: Most methods fail to generalize across diverse downstream scenarios, especially under data distribution shifts (e.g., homophilic-to-heterophilic graphs). To address these issues, we theoretically analyze existing GPL approaches and reveal that representation-level prompts essentially function as fine-tuning a simple downstream classifier, proposing that graph prompt learning should focus on unleashing the capability of pretrained models, and the classifier adapts to downstream scenarios. Based on our findings, we propose UniPrompt, a novel GPL method that adapts any pretrained models, unleashing the capability of pretrained models while preserving the structure of the input graph. Extensive experiments demonstrate that our method can effectively integrate with various pretrained models and achieve strong performance across in-domain and cross-domain scenarios.
@@ 1 Introduction
Graph Prompt Learning (GPL) GPLSurvey ; ProG , which aims to design diverse graph prompt strategies, has emerged as a promising and effective alternative paradigm that bridges between graph pretraining GCN ; GAT and downstream scenarios GFMSurvey_202506 , which overcomes the limitations of label dependency and the misalignment between upstream pretraining and downstream tasks GPPT . Most GPLs freeze the parameters of the pretrained model and tune specific prompt module. Due to its compatibility with various types of graphs, such as general graphs GPPT ; GPF , Knowledge Graphs (KGs) KGPrompt ; KGLLM ; KnowledgeGraphSurvey and Text-Attribute Graphs (TAGs) TAGSurvey ; ZeroG , GPL demonstrates strong universality and transferability, thereby improving the fields of few/zero-shot graph learning ZeroG ; TEAGLM , unifying downstream tasks AllInOne ; GraphPrompt+ ; OFA , cross-domain graph learning CrossDomainGraphLearning ; GraphLoRA , and promoting the development of Graph Foundation Models (GFMs) GFMSurvey ; GraphCLIP ; GFMTransfer .
Most GPLs can be divided into three categories according to how the prompts are integrated into the pretrained models, as illustrated in Figure 1. Input-level GPLs (Figure 1a), like feature prompt GPF and edge prompt GraphControl ; EdgePrompt , insert prompt modules or soft prompt before the pretrained models, effectively modifying the input graph to align with upstream distributions. Representation-level prompts (Figure 1c), including task tokens GPPT and prototypical subgraphs GraphPrompt ; ProNoG , applies prompts to the representations generated by the pretrained models, formulating downstream tasks that align with the pretrain objectives. Layer-wise prompt (Figure 1b) GraphPrompt+ ; MultiGPrompt combines the prompt with each layer inside the pretrained model, learning the distribution and propagation patterns in each layer. In addition, some works MultiGPrompt ; MDGFM also explore hybrid strategies that place various prompts across different layers or components (input-, representation-level) of the model.
@@
Despite the success of these explorations, our observations indicate that GPLs often experience performance instability or even negative optimization, a phenomenon also reported in recent studies ProG . Moreover, the prompt mechanisms are complex and different. There remains a lack of clear understanding of why graph prompt learning works. Through our analysis of existing methods, we identify two major issues: 1). Lack of consensus on underlying mechanisms: Although various graph prompt strategies have advanced the field, there remains no unified understanding of how these prompts interact with pretrained models. As shown in Figure 1, diverse mechanisms such as input-level, representation-level, and layer-wise prompts achieve promising performance. However, they influence the models in different ways, and the underlying interaction mechanisms are still unclear. 2). Limited scenario adaptability: Most GPLs struggle to achieve good performance on different pretrained models even in the in-domain setting. As shown in Figure 2, only fine-tuning a classifier can achieve or exceed the performance of existing GPLs. In addition, these methods have difficulty achieving excellent performance in a variety of downstream scenarios, especially when the data domains of upstream and downstream scenarios are different (e.g., from a homogeneous pretraining graph to a heterophilic downstream graph). To summarize: From the prompting mechanism to the downstream scenario, existing graph prompt learning methods exhibit an adaptation gap.
@@
To investigate the underlying mechanisms of GPL, we conduct a motivation experiment and find that existing representation-level prompt GPLs fail to consistently adapt well to different pretrained models. Moreover, they show no significant performance improvement compared to linear probe (only fine-tune the classifier), which achieves good and stable results. This motivates us to explore the relationship between different types of prompts and linear probe. Through theoretical analysis and discussions, we demonstrate that the representation-level prompt is essentially equivalent to linear probe. This primarily serves to adapt the pretrained model to downstream tasks, which focuses on fitting the outputs of the pretrained model to the downstream labels, failing to use the unique benefits of prompts. As for layer-wise methods, their reliance on layer-wise representations of the pretrained model, combined with their design complexity, makes them unsuitable. In contrast, input-level prompts avoid the limitations and preserve the advantages of prompting, they are the promising among the three categories. Therefore, we propose a perspective: graph prompt learning should focus on unleashing the capability of pretrained models, and the classifier adapts to downstream scenarios.
@@
Based on our perspective, we propose UniPrompt, a novel GPL method that adapts any pretrained models, which leverages prompt graph while preserving the original structure to unleash the capability of pretrained models. Specifically, we construct a kkNN graph as the initial prompt graph and adaptively optimize edge weights to guide message passing across nodes. To preserve the structure of the input graph, we introduce a bootstrapping strategy that integrates the prompt graph into the original graph topology, preventing model collapse and overfitting. Our main contributions can be summarized as follows:
@@ 2 Notations and Preliminary
General Graphs. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathcal{Y})$, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $|\mathcal{V}|=N$, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. These nodes are associated with feature matrix $\bm{X}\in\mathbb{R}^{N\times F}$, $\bm{X}_{i}\in\mathbb{R}^{F}$ is the feature of $v_{i}$. The edges can be represented by adjacency matrix $\bm{A}\in\{0,1\}^{N\times N}$, and $\bm{A}_{ij}=1$ iff $(v_{i},v_{j})\in\mathcal{E}$. Each node $v_{i}$ is associated with a label $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}$ denotes the set of all possible class labels. We use $P(\cdot)$ to denote the probability distribution, primarily for distinguishing concepts rather than performing mathematical derivation.
Fine-Tuning. In the pretrain-finetune paradigm, given a pretrained graph encoder $f_{\theta}$ and a downstream trainable projection head $g_{\phi}$, both parameters $\theta$ and $\phi$ are jointly optimized on a downstream dataset $\mathcal{D}=\{(\bm{A},\bm{X},y)\}$. The objective is to maximize the log-likelihood of label predictions, which can be formulated as:
$\max_{\theta,\phi}\frac{1}{|\mathcal{D}|}\sum_{(\bm{A},\bm{X},y)\in\mathcal{D}}\sum_{i=1}^{N}\log P\left(y_{i}\mid g_{\phi}\left(f_{\theta}(\bm{A},\bm{X})_{i}\right)\right),$ (1)
where $f_{\theta}(\bm{A},\bm{X})_{i}$ denotes the node representation of the node $v_{i}$. As a special case, when freezing $\theta$ (i.e., restricting $\max_{\phi}$), this reduces to linear probing where only the projection head $g_{\phi}$ is adapted.
@@
where $\mathcal{D}=\{(\mathbf{A},\mathbf{X},y)\}$ is downstream dataset, $\Psi$ represents all trainable prompt parameters, $f_{\theta}$ is the frozen pretrained encoder. $\text{Predict}_{\Psi}(\cdot)$ is a unified prediction function that takes the input graph $(\mathbf{A},\mathbf{X})$, node $v_{i}$, the pretrained encoder $f_{\theta}$, to generate the final prediction for node $v_{i}$.
For input-level prompt, $\Psi$ acts on the input $(\mathbf{A},\mathbf{X})$, transforming it before it enters $f_{\theta}$. For layer-wise prompt, $\Psi$ is embedded within the layers of $f_{\theta}$. For representation-level prompt, $\Psi$ operates on the representations from $f_{\theta}$, often as part of the classifier, directly influencing the classification.
@@ 3 Motivation Experiments and Analysis
To investigate whether graph prompt learning suffers from adaptation bias, we conduct the following experiment. As illustrated in the Figure 2, we select classic GPL methods, GPPT GPPT and GraphPrompt GraphPrompt , and compare them with fine-tuning that only optimizes the classifier. We employ widely used pretrained models: DGI DGI , GRACE GRACE , and GraphMAE GraphMAE , and we keep all other parameters consistent, and observe the differences in downstream training between prompt learning and fine-tuning. Our findings reveal that GPPT exhibits significant incompatibility when switching pretrained models, whereas GraphPrompt maintains stable performance across different pretrained models. Another observation is that even though GraphPrompt shows good convergence trends, its performance still falls short of the simple fine-tuning approach, which remains robust across different pretrained models and even outperforms GPL methods in some cases.
@@ 4 Mechanism Relationship between Prompts and Classifier
Discussion 1. Theorem 4.1 demonstrates that the representation-level prompt is fundamentally equivalent to linear probe, only designing a simple classifier can yield satisfactory results. Therefore, we suggest that graph prompt learning should focus on unleashing the capability of pretrained models, rather than adapting pretrained models to specific downstream scenarios. Specifically, an effective GPL approach should aim to combine the advantages of both mechanisms. Let $\phi_{\text{pre}}(\cdot;\mathcal{G})=\bm{A}\bm{X}\bm{W}_{\text{pre}}$ be a pretrained GNN encoder with fixed parameters $\bm{W}_{\text{pre}}$. We define $\hat{\mathcal{G}}=(\hat{\bm{A}},\hat{\bm{X}})=T(\mathcal{G})$, which modifies the input graph to align the $\phi_{\text{pre}}$. $C(\bm{h})=\bm{W}_{C}^{\top}\bm{h}$ adapts representations to downstream labels. Then, the joint optimization of $(\hat{\mathcal{G}},\bm{W}_{C})$ as follows:
$\min_{\bm{W}_{C},\hat{\mathcal{G}}}\mathcal{L}_{D}=-\sum_{v_{i}\in\mathcal{V}_{L}}y_{i}\log\sigma\left(\bm{W}_{C}^{\top}\hat{\bm{A}}\hat{\bm{X}}\bm{W}_{\text{pre}}\right)_{i},$ (4)
where $\mathcal{L}_{D}$ is the downstream task loss, typically cross-entropy loss. $\hat{\bm{A}}\hat{\bm{X}}\bm{W}_{\text{pre}}$ ensures adaptation with pretrained knowledge, and $\bm{W}_{C}$ minimizes $\mathcal{L}_{D}$ to adapt downstream task.
@@ 5 Methodology
$\min_{\phi,\Psi}\frac{1}{|\mathcal{V}_{L}|}\sum_{v_{i}\in\mathcal{V}_{L}}\ell_{D}\left(g_{\phi}\left(f_{\theta}\left(p_{\Psi}(\bm{A},\bm{X})\right)_{i}\right),y_{i}\right),$ (8)
@@ 7 Conclusion
In this work, we identify a key problem in existing Graph Prompt Learning (GPL) methods: the adaptation gap between upstream pretraining and downstream scenarios. We decompose this issue into two aspects: lack of consensus on underlying mechanisms, and limited scenario adaptability. Through motivation experiments and theoretical analysis, we reveal that the representation-level prompt is fundamentally equivalent to fine-tuning a simple downstream classifier. This primarily serves to adapt the pretrained model to downstream tasks, rather than unleashing its inherent capabilities. We propose a perspective: graph prompt learning should focus on unleashing the capability of pretrained models, and the classifier adapts to downstream scenarios. Based on our perspective, we propose UniPrompt, a novel GPL method that adapts any pretrained models, which leverages prompt-generated topology while preserving the original structure to unleash the capability of pretrained models. We evaluate UniPrompt on a comprehensive set of datasets, including homophilic and heterophilic graphs, under few-shot learning settings. The results demonstrate that UniPrompt consistently outperforms state-of-the-art baselines in both in-domain and cross-domain scenarios. Overall, our work provides new perspectives on the design principles of GPL, improving the fields of few/zero-shot graph learning, unifying downstream graph tasks, cross-domain graph learning, and promoting the development of Graph Foundation Models (GFMs).

### v2

@@ Abstract
Graph Prompt Learning (GPL) has emerged as a promising paradigm that bridges graph pretraining models and downstream scenarios, mitigating label dependency and the misalignment between upstream pretraining and downstream tasks. Although existing GPL studies explore various prompt strategies, their effectiveness and underlying principles remain unclear. We identify two critical limitations: (1) Lack of consensus on underlying mechanisms: Despite current GPLs have advanced the field, there is no consensus on how prompts interact with pretrained models, as different strategies intervene at varying spaces within the model, i.e., input-level, layer-wise, and representation-level prompts. (2) Limited scenario adaptability: Most methods fail to generalize across diverse downstream scenarios, especially under data distribution shifts (e.g., homophilic-to-heterophilic graphs). To address these issues, we theoretically analyze existing GPL approaches and reveal that representation-level prompts essentially function as fine-tuning a simple downstream classifier, proposing that graph prompt learning should focus on unleashing the capability of pretrained models, and the classifier should adapt to downstream scenarios. Based on our findings, we propose UniPrompt, a novel GPL method that adapts any pretrained models, unleashing the capability of pretrained models while preserving the input graph. Extensive experiments demonstrate that our method can effectively integrate with various pretrained models and achieve strong performance across in-domain and cross-domain scenarios.
@@ 1 Introduction
Graph Prompt Learning (GPL) GPLSurvey ; ProG , which aims to design diverse graph prompt strategies, has emerged as a promising and effective alternative paradigm that bridges between graph pretraining GCN ; GAT and downstream scenarios GFMSurvey_202506 , overcoming the limitations of label dependency and the misalignment between upstream pretraining and downstream tasks GPPT . Most GPLs freeze the parameters of the pretrained model and tune specific prompt module. Due to its compatibility with various types of graphs, such as general graphs GPPT ; GPF , Knowledge Graphs (KGs) KGPrompt ; KGLLM ; KnowledgeGraphSurvey and Text-Attribute Graphs (TAGs) TAGSurvey ; ZeroG , GPL demonstrates strong universality and transferability, thereby improving the fields of few/zero-shot graph learning ZeroG ; TEAGLM , unified task learning AllInOne ; GraphPrompt+ ; OFA , cross-domain graph learning CrossDomainGraphLearning ; GraphLoRA , and promoting the development of Graph Foundation Models (GFMs) GFMSurvey ; GraphCLIP ; GFMTransfer .
Most GPLs can be divided into three categories according to how the prompts are integrated into the pretrained models, as illustrated in Figure 1. Input-level GPLs (Figure 1a), like feature prompt GPF and edge prompt GraphControl ; EdgePrompt , insert prompt modules or soft prompt before the pretrained models, effectively modifying the input graph to align with upstream distributions in pretraining. Representation-level prompts (Figure 1c), including task tokens GPPT and prototypical subgraphs GraphPrompt ; ProNoG , applies prompts to the representations generated by the pretrained models, formulating downstream tasks that align with the pretrain objectives. Layer-wise prompt (Figure 1b) GraphPrompt+ ; MultiGPrompt combines the prompt with each layer inside the pretrained model, learning the distribution and propagation patterns in each layer. In addition, some works MultiGPrompt ; MDGFM also explore hybrid strategies that place various prompts across different layers or components (input-, representation-level) of the model.
@@
Despite the success of these explorations, our observations indicate that GPLs often experience performance instability or even negative optimization, which also be reported in recent studies ProG . Moreover, the existing prompt methods are complex and diverse. There remains a lack of clear understanding of why graph prompt learning works. Through our analysis of existing methods, we identify two major issues: 1). Lack of consensus on underlying mechanisms: Although various graph prompt strategies have advanced the field, there remains rare unified understanding of how these prompts interact with pretrained models. As shown in Figure 1, diverse mechanisms such as input-level, representation-level, and layer-wise prompts achieve promising performance. However, they influence the models in different ways, and the underlying interaction mechanisms are still unclear. 2). Limited scenario adaptability: Most GPLs struggle to achieve good performance on different pretrained models even in the in-domain setting. As shown in Figure 2, only fine-tuning a classifier can achieve or exceed the performance of existing GPLs. In addition, these methods have difficulty achieving excellent performance in a variety of downstream scenarios, especially when the data domains of upstream and downstream scenarios are different (e.g., from a homophilic pretraining graph to a heterophilic downstream graph). To summarize: From the prompting mechanism to the downstream scenario, existing graph prompt learning methods exhibit an adaptation gap.
@@
To investigate the underlying mechanisms of GPL, we conduct a motivation experiment and find that existing representation-level prompt GPLs fail to consistently adapt well to different pretrained models. Moreover, they show no significant performance improvement compared to linear probe (only fine-tune a classifier), which achieves good and stable results. This motivates us to explore the relationship between different types of prompts and linear probe. Through theoretical analysis and discussions, we demonstrate that the representation-level prompt is essentially equivalent to linear probe. This primarily serves to adapt the pretrained model to downstream tasks, which focuses on fitting the outputs of the pretrained models to the downstream labels, struggling to leverage the unique benefits of prompts. As for layer-wise methods, their reliance on layer-wise representations of the pretrained model, combined with their design complexity, makes them unsuitable. In contrast, input-level prompts avoid the limitations and preserve the advantages of prompting, they are the promising among the three categories. Therefore, we propose a perspective: graph prompt learning should focus on unleashing the capability of pretrained models, and the classifier adapts to downstream scenarios.
@@
Based on our perspective, we propose UniPrompt, a novel GPL method that adapts any pretrained models, leveraging prompt graph while preserving the original structure to unleash the capability of pretrained models. Specifically, we construct a kkNN graph as the initial prompt graph and adaptively optimize edge weights to guide message passing across nodes. To preserve the input graph, we introduce a bootstrapping strategy that integrates the prompt graph into the original graph topology, preventing model collapse and overfitting. Our main contributions can be summarized as follows:
@@ 2 Notations and Preliminary
General Graphs. Given a graph $\mathcal{G}=(\mathcal{V},\mathcal{E},\mathbf{X},\mathcal{Y})$, where $\mathcal{V}=\{v_{1},v_{2},\cdots,v_{N}\}$ is the set of nodes, $N=|\mathcal{V}|$, and $\mathcal{E}\subseteq\mathcal{V}\times\mathcal{V}$ is the set of edges. These nodes are associated with feature matrix $\mathbf{X}\in\mathbb{R}^{N\times F}$, $\mathbf{X}_{i}\in\mathbb{R}^{F}$ is the feature of $v_{i}$. The edges can be represented by adjacency matrix $\mathbf{A}\in\{0,1\}^{N\times N}$, and $\mathbf{A}_{ij}=1$ iff $(v_{i},v_{j})\in\mathcal{E}$. Each node $v_{i}$ is associated with a label $y_{i}\in\mathcal{Y}$, where $\mathcal{Y}$ denotes the set of all possible class labels. We use $P(\cdot)$ to denote the probability distribution, primarily for distinguishing concepts rather than performing mathematical derivation.
Fine-Tuning. In the pretrain-finetune paradigm, given a pretrained graph encoder $f_{\theta}$ and a downstream trainable projection head $g_{\phi}$, both parameters $\theta$ and $\phi$ are jointly optimized on a downstream dataset $\mathcal{D}=\{(\mathbf{A},\mathbf{X},y)\}$. The objective is to maximize the log-likelihood of label predictions, which can be formulated as:
$\max_{\theta,\phi}\frac{1}{|\mathcal{D}|}\sum_{(\mathbf{A},\mathbf{X},y)\in\mathcal{D}}\sum_{i=1}^{N}\log P\left(y_{i}\mid g_{\phi}\left(f_{\theta}(\mathbf{A},\mathbf{X})_{i}\right)\right),$ (1)
where $f_{\theta}(\mathbf{A},\mathbf{X})_{i}$ denotes the node representation of the node $v_{i}$. As a special case, when freezing $\theta$ (i.e., restricting $\max_{\phi}$), this reduces to linear probing where only the projection head $g_{\phi}$ is adapted.
@@
where $\mathcal{D}=\{(\mathbf{A},\mathbf{X},y)\}$ is a downstream dataset, $\Psi$ represents all trainable prompt parameters, $f_{\theta}$ is the frozen pretrained encoder. $\text{Predict}_{\Psi}(\cdot)$ is a unified prediction function that takes the input graph $(\mathbf{A},\mathbf{X})$, node $v_{i}$, the pretrained encoder $f_{\theta}$, to predict the label of node $v_{i}$.
For input-level prompt, $\Psi$ acts on the input $(\mathbf{A},\mathbf{X})$, transforming it before it enters $f_{\theta}$. For layer-wise prompt, $\Psi$ is integrated within the layers of $f_{\theta}$. For representation-level prompt, $\Psi$ operates on the representations from $f_{\theta}$, often as part of the classifier, directly influencing the downstream tasks.
@@ 3 Motivation Experiments and Analysis
To investigate whether graph prompt learning suffers from adaptation bias, we conduct the following experiment. As illustrated in the Figure 2, we select classic GPL methods, GPPT GPPT and GraphPrompt GraphPrompt , and compare them with fine-tuning that only optimizes the classifier. We employ widely used pretrained models: DGI DGI , GRACE GRACE , and GraphMAE GraphMAE , and we keep all other parameters consistent, and observe the differences in downstream training between prompt learning and fine-tuning. Our experimental results reveal that GPPT exhibits significant incompatibility when switching pretrained models, whereas GraphPrompt maintains stable performance across different pretrained models. Another observation is that even though GraphPrompt shows good convergence trends, its performance still falls short of the simple fine-tuning approach in many scenarios, which remains robust across different pretrained models and even outperforms GPL methods in some cases.
@@ 4 Mechanism Relationship between Prompts and Classifier
Discussion 1. Theorem 4.1 demonstrates that the representation-level prompt is fundamentally equivalent to linear probe, only designing a simple classifier can yield satisfactory results. Therefore, we suggest that graph prompt learning should focus on unleashing the capability of pretrained models, rather than adapting pretrained models to specific downstream scenarios. Specifically, an effective GPL approach should aim to combine the advantages of both mechanisms. Let $\phi_{\text{pre}}(\cdot;\mathcal{G})=\mathbf{A}\mathbf{X}\mathbf{W}_{\text{pre}}$ be a pretrained GNN encoder with fixed parameters $\mathbf{W}_{\text{pre}}$. We define $\hat{\mathcal{G}}=(\hat{\mathbf{A}},\hat{\mathbf{X}})=T(\mathcal{G})$, which modifies the input graph to align the $\phi_{\text{pre}}$. $C(\mathbf{h})=\mathbf{W}_{C}^{\top}\mathbf{h}$ adapts representations to downstream labels. Then, the joint optimization of $(\hat{\mathcal{G}},\mathbf{W}_{C})$ as follows:
$\min_{\mathbf{W}_{C},\hat{\mathcal{G}}}\mathcal{L}_{D}=-\sum_{v_{i}\in\mathcal{V}_{L}}y_{i}\log\sigma\left(\mathbf{W}_{C}^{\top}\hat{\mathbf{A}}\hat{\mathbf{X}}\mathbf{W}_{\text{pre}}\right)_{i},$ (4)
where $\mathcal{L}_{D}$ is the downstream task loss, typically cross-entropy loss. $\hat{\mathbf{A}}\hat{\mathbf{X}}\mathbf{W}_{\text{pre}}$ ensures adaptation with pretrained knowledge, and $\mathbf{W}_{C}$ minimizes $\mathcal{L}_{D}$ to adapt downstream task.
@@ 5 Methodology
$\min_{\phi,\Psi}\frac{1}{|\mathcal{V}_{L}|}\sum_{v_{i}\in\mathcal{V}_{L}}\ell_{D}\left(g_{\phi}\left(f_{\theta}\left(p_{\Psi}(\mathbf{A},\mathbf{X})\right)_{i}\right),y_{i}\right),$ (8)
@@ 7 Conclusion
In this work, we categorize existing Graph Prompt Learning (GPL) methods based on their mechanisms and conduct an analysis of them. Through this analysis, we identify a key problem in existing GPL methods: the adaptation gap between upstream pretraining and downstream scenarios. We decompose this issue into two aspects: lack of consensus on underlying mechanisms, and limited scenario adaptability. Through motivation experiments and theoretical analysis, we reveal that the representation-level prompt is fundamentally equivalent to fine-tuning a simple downstream classifier. This primarily serves to adapt the pretrained model to downstream tasks, rather than unleashing its inherent capabilities. We propose a perspective: graph prompt learning should focus on unleashing the capability of pretrained models, and the classifier adapts to downstream scenarios. Based on our perspective, we propose UniPrompt, a novel GPL method that adapts any pretrained models, which leverages prompt-generated topology while preserving the original structure to unleash the capability of pretrained models. We evaluate UniPrompt on a comprehensive set of datasets, including homophilic and heterophilic graphs, under few-shot learning settings. The results demonstrate that UniPrompt consistently outperforms state-of-the-art baselines in both in-domain and cross-domain scenarios. Overall, our work provides new perspectives on the design principles of GPL, improving the fields of few/zero-shot graph learning, unifying downstream graph tasks, cross-domain graph learning, and promoting the development of Graph Foundation Models.